# FedLog: Personalized Federated Classification with Less Communication and More Flexibility

### v1

@@ 3.1 FedLog
     $\boldsymbol{\eta}\leftarrow\operatorname*{arg\,max}_{\boldsymbol{\eta}}\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)$ (Eq.  15)
Based on the current model, we propose a new algorithm named FedLog, as shown in Alg. 1. At the beginning, the server initializes $\boldsymbol{\eta}$ (the global head) randomly, and the clients initializes $\tilde{\boldsymbol{\theta}}_{c}$ (the local bodies) either completely randomly, or with the same random seed sent by the server to unify the initialization. Note $\tilde{\boldsymbol{\theta}}_{c}$ is not part of our exponential family assumption, thus we do not require them to have the same shape or architecture amongst different clients.
$\begin{split}\mathcal{L}_{c}&=-\sum_{i=1}^{n_{c}}\ln\Pr({\bf y}_{c,i}|\boldsymbol{\Phi}_{c,i},\boldsymbol{\eta})=-\sum_{i=1}^{n_{c}}\ln\frac{\exp(\boldsymbol{\eta}_{y_{c,i}}^{\top}\boldsymbol{\Phi}_{c,i})}{\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\Phi}_{c,i})}\end{split}$ (13)
where $(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})\in\mathbb{R}^{m}\times\mathbb{R}$ is the $i^{th}$ data point of $\boldsymbol{\Phi}_{c},{\bf y}_{c}$. This is exactly the cross entropy loss widely used in deep learning for classification tasks. Then, clients compute sufficient statistics of their local data $\sum_{i=1}^{n_{c}}{\bf T}(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})=\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}$. They then send the summations and $n_{c}$ to the server for global head learning. As we have discussed in Sec. 2.2, the sufficient statistics contain all the information in the raw dataset that could be used to infer $\boldsymbol{\eta}$ in our model. In this case, the server only needs to know the summation of all the sufficient statistics, since
$\begin{split}&\Pr(\boldsymbol{\eta}|\boldsymbol{\Phi}_{c_{1}},{\bf y}_{c_{1}},\cdots,\boldsymbol{\Phi}_{c_{k}},{\bf y}_{c_{k}})=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},\nu+\sum_{c\in S}n_{c})\end{split}$ (14)
$\begin{split}\boldsymbol{\eta}&=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\frac{\exp\left(\boldsymbol{\eta}^{\top}(\boldsymbol{\chi}+\boldsymbol{\Phi})\right)}{(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{(\nu+n)}}\end{split}$ (15)
@@
Note Eq. 15 is a convex optimization task, and we can easily compute its sole maximum by gradient descent. Then, the server sends $\boldsymbol{\eta}$ back to all the clients and starts the next round of updates. The process is repeated until convergence.
3.2 Interpret FedLog
@@
In this section, we analyze our assumptions in more details and give some insights about how FedLog works. The first assumption we made is that $\forall c\in S$, the local data points $\boldsymbol{\Phi}_{c},{\bf y}_{c}$ are from the same exponential family distribution whose pdf is given by Eq. 9. In other words, we assume that the local bodies $\tilde{\boldsymbol{\theta}}_{c}$ transform their input, of any form, to the same representation space. This may seem infeasible at the first glance, since we do not directly aggregate the local body parameters, and they may take any architecture. However, note we fix the global head $\boldsymbol{\eta}$ during local updates, by which the local bodies are forced to learn a universal representation space. See Sec. 4.1 for more details. This assumption allows principled discriminative training for the local bodies with cross entropy loss, but unavoidably leaves a generative model for learning the global head. We can further see that $\Pr(\boldsymbol{\phi}|y,\boldsymbol{\eta})\propto\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}-\boldsymbol{\phi}^{\top}\boldsymbol{\phi})$, which is the kernel of a multivariate Gaussian distribution. This means we essentially assumed a mixture of Gaussians for the local features $\boldsymbol{\phi}$. This is a mild assumption, since sufficiently expressive DNN encoders can map data to any distribution any closely.
@@
The second assumption we made is the prior of the global head. Since $\boldsymbol{\eta}$ has a support over $\mathbb{R}^{m*n_{class}}$, it is impossible to specify a uniform prior. Without any prior knowledge, we can set ${\boldsymbol{\chi}}={\bf 0},\nu=1$. The prior then becomes $\Pr(\boldsymbol{\eta})\propto\exp(-A(\boldsymbol{\eta}))=(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{-1}$. The pdf takes its maximum at $\boldsymbol{\eta}={\bf 0}$ and decreases quickly as the absolute value of entries of $\boldsymbol{\eta}_{y}$ grows larger. This is in analogy to the Lasso regularizer in the regression case, which prevents the model from learning coefficients with large absolute values that are more likely due to noise or over-fitting.
@@
From the Bayesian view, FedLog can take any deep classifier, and make the last linear layer Bayesian. It essentially operates a Bayesian logistic regression model on the local representations, thus the name FedLog. We start from a generative assumption and achieve the cross entropy loss conditional likelihood usually assumed directly in Bayesian logistic regression. We obtained an analytical solution for the kernel of the posterior, which can be calculated easily by the summation of sufficient statistics, whose sizes are as small as the size of the last linear layer, and avoid the need to share separated data points.
From the federated representation learning view, FedLog has a one-layer global head and a deep local body. It iterates between learning local representations and learning global linear separators as if it has seen all the local representations. The two learning processes are completely separated, unlike the common paradigm where the local representations and the linear separators are often optimized jointly. The clients are only responsible for moving local representations to the correct sides of the fixed linear separator. The server is only responsible for finding the best linear separator given the local representations. As we will show with the experiments, we can automatically learn universal local representations and converge faster than using the averaging heuristic.
3.3 Differential Privacy
FL can be combined with formal mechanisms such as differential privacy (DP) [30, 31] or secure multi-party computation (MPC) [32, 33, 34], to provide formal privacy guarantees. We now extend FedLog to be differentially private.
Theorem 3.1.
If the absolute value of features are clipped to $b$, FedLog messages satisfy $(\epsilon,\delta)$-DP with additive Gaussian noise ${\bf T}^{\prime}(\boldsymbol{\Phi}_{c},{\bf y}_{c}):={\bf T}(\boldsymbol{\Phi}_{c},{\bf y}_{c})+\mathcal{N}({\bf 0},\sigma^{2}{\bf I})$, where $\sigma=\sqrt{2(1+(m-1)*b^{2})\ln(1.25/\delta)}/{\epsilon}$.
@@ 3.3 Differential Privacy
$L_{2}({\bf T})=\max_{\mathcal{D},\mathcal{D}^{\prime}}||{\bf T}(\mathcal{D})-{\bf T}(\mathcal{D}^{\prime})||_{2}=\max_{\boldsymbol{\phi},y}||{\bf T}(\boldsymbol{\phi},y)||_{2}=\sqrt{1+(m-1)*b^{2}}$
See more details in Appendix A. ∎
When $n$ is large enough, the amount of noise (independent of $n$) becomes negligible to the model. Optionally, with secure MPC, it is sufficient to add such Gaussian noise once globally to the summation of all the messages each global update round. The server then computes the posterior normally with the noisy sufficient statistics.
4 Experiments
4.1 Synthetic
Table 1: Testing accuracy and communication cost reported for MNIST, CIFAR10, and CIFAR100. Accuracy reports the mean $\pm$ standard error of the testing accuracy over $10$ seeds. Higher is better. Communication cost reports the number of bits transmitted between one client and the server each global update round. Lower is better. $\Uparrow$ denotes significantly better results with $p<0.01$; $\Downarrow$ denotes significantly worse results with $p<0.01$.
@@ 4.1 Synthetic
FedAvg 89.76$\pm$0.69 698880 26.29$\pm$0.44 4102528 13.34$\pm$0.15 4684288
LG-FedAvg 1 97.85$\pm$0.05 16320 86.57$\pm$0.29 64640 55.00$\pm$0.26 646400
LG-FedAvg 2 98.18$\pm$0.06 529920 85.56$\pm$0.32 839040 54.90$\pm$0.24 1420800
FedPer 96.16$\pm$0.19 168960 83.54$\pm$0.40 3263488 52.82$\pm$0.21 3263488
FedRep 95.51$\pm$0.29 168960 82.96$\pm$0.35 3263488 48.70$\pm$0.29 3263488
CS-FL 79.65$\pm$1.22 72088 23.60$\pm$1.08 423092 4.52$\pm$0.15 966172
FedBabu 86.30$\pm$1.04 682560 25.37$\pm$0.44 4037888 9.70$\pm$0.16 4037888
FedProto 98.19$\pm$0.06 16320 87.37$\pm$0.26 64640 55.32$\pm$0.19 646400
FedDBE 96.79$\pm$0.34 700480 72.77$\pm$0.79 4108928 36.67$\pm$0.85 4690688
FedLog (ours) 98.15$\pm$0.05 16320 87.08$\pm$0.22 64640 56.46$\pm$0.27 646400
Table 2: Lower is better. $\Uparrow$ denotes significantly worse results with $p<0.01$.
MNIST CIFAR10 CIFAR100
rounds needed rounds needed rounds needed
LG-FedAvg 1 5.9$\pm$0.38 40.3$\pm$3.09 67.1$\pm$2.76
FedProto 3.7$\pm$0.21 28.5$\pm$2.05 46.6$\pm$1.90
FedLog(ours) 3.9$\pm$0.38 21.5$\pm$2.23 36.8$\pm$1.39
4.2 Communication Cost
@@
We conduct experiments on MNIST, CIFAR10, and CIFAR100, to show FedLog achieves better accuracy with less communication. We compare FedLog with the following baselines: i) FedAvg [2], which averages the whole model each global update round; ii) LG-FedAvg [3], which localizes bodies and averages heads. Two variants are reported: LG-FedAvg 1 which maintains one global layer and LG-FedAvg 2 which maintains two global layers; iii) FedPer [4], which localizes heads and averages bodies; iv) FedRep [5], which also localizes heads and averages bodies, but trains local heads and bodies separately; v) CS-FL [24], a model compression technique that compresses messages with the compressed sensing framework; vi) FedBabu [35], which averages bodies and never updates heads; vii) FedProto [12], which averages local feature representations by class and forces local models to learn similar representations; viii) FedDBE [36], which averages both bodies and heads but accelerates convergence by learning a domain representation bias. Convolutional neural networks of the same architecture and initialization is used for all the algorithms. FedPer and FedRep localizes the last two fully connected layers for less communication cost.
@@
We first distribute the training set evenly into $50$ (for MNIST) or $100$ (for CIFAR10 and CIFAR100) clients. Each client takes only $2$ classes to maximize the heterogeneity. The testing set is distributed similarly to the clients, following the same distribution as the training data. Specially, for MNIST only, to simulate the difficult situation where clients do not have sufficient local data, we train on only $5\%$ of the training set, while test on the whole testing set. All the clients start from the same initialization, and all the algorithms are run for $100$ (for MNIST and CIFAR10) or $150$ (for CIFAR100) global update rounds. The testing accuracy is recorded after each round. Hyperparameters are optimized beforehand through grid search. See Appendix B for the model architectures, hyperparameters used by each algorithm and other experiment details.
@@
We report the mean and the standard error of the testing accuracy resulting from $10$ seeds in Table 1. We measured the statistical significance of the results compared to FedLog with one-tailed Wilcoxon signed-rank tests [37]. We also report the communication cost of each global update round, namely the number of bits transmitted between one client and the server. The results show FedLog can achieve statistically significant accuracy improvement or at least comparable utility compared to prior arts, with the least communication cost as small as $2.3\%$, $1.6\%$, and $13.8\%$ of the communication cost of FedAvg for MNIST, CIFAR10, and CIFAR100 respectively.
We further report how many global rounds are needed to reach a reasonable accuracy threshold in Table 2. The thresholds are 97%, 83%, and 53% respectively. These numbers are the largest integer of accuracy which all three algorithms have reached in 10 seeds. The results show FedLog needs less rounds to converge than LG-FedAvg and FedProto, thus saving more communication cost despite the cost per round is the same.
4.3 Celeba
We compare FedLog to LG-FedAvg and FedProto on large dataset Celeba [38] preprocessed by LEAF [39]. We sampled 70838 images from 2360 clients. Each client represents a celebrity, and the local data are images of the same person. Sample inputs are shown in the left half of Fig. 2. The task is to classify if the celebrity is smiling. MobileNetV2 [40] is implemented as the classifier model for all 3 algorithms. To test how these algorithms perform under the situation where only a few communication rounds are available, we run every algorithm to optimum locally before global aggregation. The results are shown in the right half of Fig. 2. FedLog performs statistically significantly better than FedProto (71.19$\pm$0.45 v.s. 69.98$\pm$0.16, $p=0.03$) and LG-FedAvg (66.84$\pm$0.19, $p=0.001$).
4.4 Flexible Architecture
We simulate the situation where clients have different computational resources on CIFAR100. We randomly select half of the clients and assign them a smaller convolutional neural network, where the second last fully connected layer is removed. Different clients start from different local body initialization, but the global head is unified. Other experiment details are similar to Sec. 4.2. We compare FedLog to LG-FedAvg 1 and FedProto. As shown in the left graph of Fig. 3, FedLog converges faster than LG-FedAvg and FedProto. The accuracy of FedLog is also statistically significantly higher than the accuracy of FedProto (55.31$\pm$0.15 v.s. 54.74$\pm$0.13, $p=0.002$) and LG-FedAvg (54.03$\pm$0.08, $p=0.001$).
Figure 2: Left: sample Celeba input. Right: model accuracy. Colored area shows mean $\pm$ standard deviation of the accuracy.
@@ 4.4 Flexible Architecture
Figure 3: Left: flexible architecture results. Right: differential privacy trade-off. Colored area shows mean $\pm$ standard deviation of the accuracy.
4.5 Differential Privacy
@@
We conduct experiments to show the trade-off between privacy budget $\epsilon$ and the accuracy of FedLog on CIFAR10. Experiment details are similar to Sec. 4.2, except we add an activation function to clip the extracted features to $b=2$, and $\delta=0.01$. As shown in the right graph of Fig. 3, the accuracy of differentially private FedLog quickly grows back to optimum when $\epsilon\geq 0.1$. This is a very strong privacy budget, and the results show FedLog performs well without sacrificing clients’ privacy.
5 Conclusion

### v2

@@ Introduction
     $\boldsymbol{\eta}\leftarrow\operatorname*{arg\,max}_{\boldsymbol{\eta}}\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)$ (Eq.  14)
@@
Note Eq. 14 is a convex optimization task. We can easily compute its sole maximum by gradient descent with complexity $O(m*n_{class})$. Then, the server sends $\boldsymbol{\eta}$ back to all the clients and starts the next round of updates. The process is repeated until convergence.
Interpretation and FedLog-C
@@
In this section, we analyze our assumptions in more details and give some insights about how FedLog works. The first assumption we made is that $\forall c\in S$, the local data points $\boldsymbol{\Phi}_{c},{\bf y}_{c}$ are from the same exponential family distribution whose pdf is given by Eq. 4. In other words, we assume that the local bodies $\tilde{\boldsymbol{\theta}}_{c}$ transform their input, of any form, to the same representation space. This may seem infeasible at first glance since we do not directly aggregate the local body parameters, and they may follow any architecture. However, note we fix the global head $\boldsymbol{\eta}$ during local updates, by which the local bodies are forced to learn a universal representation space. See the synthetic experiment below for more details. This assumption allows principled discriminative training for the local bodies with cross entropy loss, but unavoidably leaves a generative model for learning the global head. We can further see that $\Pr(\boldsymbol{\phi}|y,\boldsymbol{\eta})\propto\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}-\boldsymbol{\phi}^{\top}\boldsymbol{\phi})$, which is the kernel of a multivariate Gaussian distribution. This means we essentially assumed a mixture of Gaussians for the local features $\boldsymbol{\phi}$. To mitigate the gap between the assumption and the actual feature distribution, we propose a variation FedLog-C. An auxiliary loss is added during local training to force the local bodies to learn Gaussian-like clusters. Let $\boldsymbol{\Phi}_{y}$ denote the $((y-1)*m)^{th}$ to $(y*m)^{th}$ entries of $\boldsymbol{\Phi}$. Let $\boldsymbol{\Phi}_{y,0}$ denote the first entry of $\boldsymbol{\Phi}_{y}$. Then $\overline{\boldsymbol{\Phi}_{y}}=\boldsymbol{\Phi}_{y}/\boldsymbol{\Phi}_{y,0}$ is the global mean representation of class $y$. Inspired by contrastive learning (Schroff, Kalenichenko, and Philbin 2015), we derive the new local loss as:
$\begin{split}\mathcal{L}^{\prime}_{c}=\mathcal{L}_{c}&+\alpha\sum_{i=1}^{n_{c}}(\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}-\overline{\boldsymbol{\Phi}_{{\bf y}_{c,i}}})^{2}/n_{c}\\ &-\beta\sum_{y^{\prime}\neq{\bf y}_{c,i}}\sum_{i=1}^{n_{c}}(\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}-\overline{\boldsymbol{\Phi}_{y^{\prime}}})^{2}/n_{c}\end{split}$ (15)
$\alpha\in\mathbb{R}_{\geq 0}$ controls how compact the clusters should be. $\beta\in\mathbb{R}_{\geq 0}$ controls the distance between different clusters. Since clients need to know $\boldsymbol{\Phi}$, the server simply broadcasts the aggregated statistic to all the clients. Clients can optimize the same $\boldsymbol{\eta}$ locally, preserving the same communication cost.
@@
The second assumption we made is the prior of the global head. Since $\boldsymbol{\eta}$ has a support over $\mathbb{R}^{m*n_{class}}$, it is impossible to specify a uniform prior. Without any prior knowledge, we can set ${\boldsymbol{\chi}}={\bf 0},\nu=1$. The prior then becomes $\Pr(\boldsymbol{\eta})\propto\exp(-A(\boldsymbol{\eta}))=(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{-1}$. The p.d.f. takes its maximum at $\boldsymbol{\eta}={\bf 0}$ and decreases quickly as the absolute value of entries of $\boldsymbol{\eta}_{y}$ grows larger. This is in analogy to the Lasso regularizer in the regression case, which prevents the model from learning coefficients with large absolute values due to noise or over-fitting.
@@
From the Bayesian view, FedLog can take any deep classifier, and make the last linear layer Bayesian. It essentially operates a Bayesian logistic regression model on the local representations. We start from a generative assumption and achieve the cross-entropy loss conditional likelihood usually assumed directly in Bayesian logistic regression. We obtained an analytical solution for the kernel of the posterior, which can be calculated easily by the summation of sufficient statistics. The shared statistic cannot be further compressed without losing information from the representations. We formalize this statement with the following theorem.
Theorem 2.
If $({\boldsymbol{\phi}_{1}},y_{1}),({\boldsymbol{\phi}_{2}},y_{2}),\cdots,({\boldsymbol{\phi}_{n}},y_{n})$ are i.i.d. samples from the exponential family defined with Eq. 4, then ${\bf T}(({\boldsymbol{\phi}_{1}},y_{1}),\cdots,({\boldsymbol{\phi}_{n}},y_{n}))=\sum_{i=1}^{n}{\boldsymbol{\phi}_{i}}\otimes{\bf e}_{y_{i}}$ is a minimal sufficient statistic independent of every ancillary statistic.
@@
See Appendix B for the proof. ∎
From the federated representation learning view, FedLog has a one-layer global head and a deep local body. It iterates between learning local representations and learning global linear separators as if it has seen all the local representations. The two learning processes are completely separated, unlike the common paradigm where the local representations and the linear separators are often optimized jointly. The clients are only responsible for moving local representations to the correct sides of the fixed linear separator. The server is only responsible for finding the best linear separator given the local representations. As we will show with the experiments, we can converge faster than using the averaging heuristic.
Differential Privacy
FL can be combined with formal mechanisms such as differential privacy (DP) (Wei et al. 2020; Triastcyn and Faltings 2019) or secure multi-party computation (MPC) (Truex et al. 2019; Byrd and Polychroniadou 2020; Li et al. 2020), to provide formal privacy guarantees. We now extend FedLog to be differentially private.
($\epsilon,\delta$)-DP protects clients’ privacy by adding noise to the shared information so that the adversaries cannot effectively tell if any record is included in the dataset (controlled by $\epsilon>0$) at most times (controlled by $10\leq\delta<$ 1) (Kerkouche et al. 2021).
Definition 3.
A mechanism $M_{DP}$ satisfies $(\epsilon$, $\delta)$-DP if for any two datasets $\mathcal{D},\mathcal{D}^{\prime}$ that differ by only one record (i.e. $|(\mathcal{D}-\mathcal{D}^{\prime})\cup(\mathcal{D}^{\prime}-\mathcal{D})|=1$), and for any possible output $O\in Range(M_{DP})$,
Intuitively, $(\epsilon,\delta)$-DP guarantees that the inner log ratio, considered as the information loss leaked to the adversaries, is bounded by the privacy budget $\epsilon$ with probability $\delta$. We add Gaussian noise to shared sufficient statistics as follows.
Theorem 4.
If the absolute value of features are clipped to $b$ and there are in total $k$ global update rounds, FedLog messages satisfy $(\epsilon,\delta)$-DP with additive Gaussian noise ${\bf T}^{\prime}(\boldsymbol{\Phi}_{c},{\bf y}_{c}):={\bf T}(\boldsymbol{\Phi}_{c},{\bf y}_{c})+\mathcal{N}({\bf 0},\sigma^{2}{\bf I})$, where $\sigma=\sqrt{8k(1+(m-1)*b^{2})\ln(e+\epsilon/\delta)}/{\epsilon}$.
Proof.
$\max_{\mathcal{D},\mathcal{D}^{\prime}}||{\bf T}(\mathcal{D})-{\bf T}(\mathcal{D}^{\prime})||_{2}$ $=\max_{\boldsymbol{\phi},y}||{\bf T}(\boldsymbol{\phi},y)||_{2}$
$=\sqrt{1+(m-1)*b^{2}}$
See more details in Appendix C. ∎
When $n$ is large enough, the amount of noise (independent of $n$) becomes negligible to the model. Optionally, with secure MPC or central DP, it is sufficient to add such Gaussian noise once globally to $\boldsymbol{\Phi}$ each global update round. The server then computes the posterior normally with the noisy sufficient statistics.
Table 1: Testing accuracy and communication cost reported for MNIST, CIFAR10, and CIFAR100. Accuracy reports the mean $\pm$ standard error of the testing accuracy over $10$ seeds. Higher is better. Communication cost reports the total message size transmitted between clients and the server. Lower is better. $\Uparrow$ denotes significantly higher results with $p<0.01$; $\Downarrow$ denotes significantly lower results with $p<0.01$.
@@
FedAvg 89.76$\pm$0.69 3.45$\pm$0.02Gb 26.29$\pm$0.44 37.9$\pm$1.31Gb 13.34$\pm$0.15 69.1$\pm$0.47Gb
LG-FedAvg 1 97.85$\pm$0.05 4.81$\pm$0.31Mb 86.57$\pm$0.29 0.26$\pm$0.02Gb 55.00$\pm$0.26 4.34$\pm$0.18Gb
LG-FedAvg 2 98.18$\pm$0.06 0.16$\pm$0.07Gb 85.56$\pm$0.32 3.38$\pm$0.32Gb 54.90$\pm$0.24 9.53$\pm$0.53Gb
FedPer 96.16$\pm$0.19 0.65$\pm$0.05Gb 83.54$\pm$0.40 27.7$\pm$1.50Gb 52.82$\pm$0.21 42.7$\pm$1.47Gb
FedRep 95.51$\pm$0.29 36.3$\pm$3.38Mb 82.96$\pm$0.35 15.3$\pm$1.33Gb 48.70$\pm$0.29 11.0$\pm$0.40Gb
CS-FL 79.65$\pm$1.22 0.35$\pm$0.01Gb 23.60$\pm$1.08 2.72$\pm$0.41Gb 4.52$\pm$0.15 13.7$\pm$0.22Gb
FedBabu 86.30$\pm$1.04 2.25$\pm$0.02Gb 25.37$\pm$0.44 34.7$\pm$2.29Gb 9.70$\pm$0.16 59.3$\pm$0.31Gb
FedProto 98.19$\pm$0.06 3.02$\pm$0.17Mb 87.37$\pm$0.26 0.18$\pm$0.01Gb 55.32$\pm$0.19 3.01$\pm$0.12Gb
FedDBE 96.79$\pm$0.34 1.71$\pm$0.39Gb 72.77$\pm$0.79 38.3$\pm$0.95Gb 36.67$\pm$0.85 55.1$\pm$2.80Gb
FedLog (ours) 98.15$\pm$0.05 3.18$\pm$0.31Mb 87.08$\pm$0.22 0.14$\pm$0.01Gb 56.46$\pm$0.27 2.38$\pm$0.09Gb
FedLog-C (ours) 98.41$\pm$0.07 3.18$\pm$0.15Mb 87.57$\pm$0.25 0.11$\pm$0.01Gb 56.78$\pm$0.26 2.74$\pm$0.12Gb
Experiments
Synthetic
Communication Cost
@@
To show FedLog achieves better accuracy with less communication with non-i.i.d. clients, we conduct experiments on MNIST, CIFAR10, and CIFAR100. We compare FedLog and FedLog-C with the following baselines: i) FedAvg (McMahan et al. 2017), which averages the whole model each global update round; ii) LG-FedAvg (Liang et al. 2020), which localizes bodies and averages heads. Two variants are reported: LG-FedAvg 1 which maintains one global layer and LG-FedAvg 2 which maintains two global layers; iii) FedPer (Arivazhagan et al. 2019), which localizes heads and averages bodies; iv) FedRep (Collins et al. 2021), which also localizes heads and averages bodies, but trains local heads and bodies separately; v) CS-FL (Li, Li, and Varshney 2021), a model compression technique that compresses messages with the compressed sensing framework; vi) FedBabu (Oh, Kim, and Yun 2021), which averages bodies and never updates heads; vii) FedProto (Tan et al. 2022b), which averages local feature representations by class and forces local models to learn similar representations; viii) FedDBE (Zhang et al. 2024), which averages both bodies and heads but accelerates convergence by learning a domain representation bias. Convolutional neural networks of the same architecture and initialization is used for all the algorithms. FedPer and FedRep localizes the last two layers.
@@
We first distribute the training set into $(50,100,100)$ heterogeneous clients for (MNIST, CIFAR10, CIFAR100) respectively. Each client takes only $(2,2,10)$ classes. The testing set is distributed similarly to the clients, following the same distribution as the training data. Specially, for MNIST only, to simulate the difficult situation where clients do not have sufficient local data, we train on $5\%$ of the training set, while test on the whole testing set. All the clients start from the same initialization, and all the algorithms are run for $(100,100,150)$ global update rounds. The testing accuracy is recorded after each round. Hyperparameters are optimized beforehand through grid search. See Appendix D for the model architectures, hyperparameters used by each algorithm and other experiment details.
@@
We report mean $\pm$ standard error of the testing accuracy resulting from $10$ seeds in Table 1. We measured the statistical significance of the results compared to FedLog-C with one-tailed Wilcoxon signed-rank tests (Wilcoxon 1992). Similarly, we report the total communication cost, namely the total size of messages transmitted between clients and the server, to reach a reasonable accuracy threshold (97%, 83%, and 53% respectively). If it is never reached, we stop counting at the round with the highest accuracy. These thresholds are the largest integer of accuracy which all competitive algorithms (FedLog, LG-FedAvg, FedProto) have reached in 10 seeds.
The results show FedLog can achieve statistically significant accuracy improvement compared to prior arts, with the least communication cost as small as $0.09\%$, $0.29\%$, and $3.44\%$ of the communication cost of FedAvg for MNIST, CIFAR10, and CIFAR100 respectively. The closest baseline is FedProto, which has a slightly lower performance and requires more communication on CIFAR10 and CIFAR100 because it uses FedAvg instead of Bayesian inference for aggregation.
@@
Figure 2: From top to bottom: Celeba, flexible architecture, and differential privacy results. Colored area shows mean $\pm$ standard deviation of the accuracy.
Celeba
We compare FedLog and FedLog-C to LG-FedAvg and FedProto on the large dataset Celeba (Liu et al. 2015) preprocessed by LEAF (Caldas et al. 2018). We sampled 70838 images from 2360 clients. Each client represents a celebrity, and the local data are images of the same person. The task is to classify if the celebrity is smiling. MobileNetV2 (Sandler et al. 2018) is implemented as the classifier model for all algorithms. To test how these algorithms perform under the situation where only a few communication rounds are available, we run every algorithm to optimum locally before global aggregation. The results are shown in the top graph of Fig. 2. FedLog-C performs statistically significantly better than FedProto (71.96$\pm$0.18 v.s. 69.98$\pm$0.16, $p=0.001$), LG-FedAvg (66.84$\pm$0.19, $p=0.001$), and FedLog (71.19$\pm$0.45, $p=0.05$).
Flexible Architecture
We simulate the situation where clients have different computational resources on CIFAR100. We randomly select half of the clients and assign them a smaller convolutional neural network, where the second last fully connected layer is removed. Different clients start from different local body initialization, but the global head is unified. We compare FedLog and FedLog-C to LG-FedAvg 1 and FedProto. As shown in the middle graph of Fig. 2, FedLog-C and FedLog converges faster than LG-FedAvg and FedProto. The accuracy of FedLog-C is also statistically significantly higher than FedProto (55.86$\pm$0.11 v.s. 54.74$\pm$0.13, $p=0.001$), LG-FedAvg (54.03$\pm$0.08, $p=0.001$), and FedLog (55.31$\pm$0.15, $p=0.001$).
Differential Privacy
@@
We conduct experiments to show the trade-off between privacy budget $\epsilon$ and the accuracy of FedLog on CIFAR10. We add an activation function to clip the extracted features to $b=2$. Following a common practice in FL (Wei et al. 2020), we set $\delta=0.01$. As shown in the bottom graph of Fig. 2, the accuracy of differentially private FedLog quickly grows back to optimum when $\epsilon\geq 0.5$, if the server is trusted or MPC is implemented so central DP is applicable. This is a strong privacy budget that shows FedLog performs well without sacrificing clients’ privacy. Otherwise, local DP can have more impact on the model utility, but the accuracy is still acceptable when $\epsilon\geq 5.0$.
Limitations
One limitation of FedLog is that the algorithm works only for classification. The model is solely designed to mimic cross-entropy loss and therefore a different loss function with a different model would be needed for regression. FedLog also assumes an exponential family distribution with a prior of the form $exp(-A(\boldsymbol{\eta}))$ and local transformed features distributed according to a mixture of Gaussians. However the mixture of Gaussian assumption is mitigated in FedLog-C by introducing an auxiliary loss that helps satisfy the assumption. A benefit of the exponential family and mixture of Gaussian assumptions is that the data summaries are provable sufficient statistics (Thm. 2) and we obtain a closed form solution for the normalization constant (Eq. 11).
Conclusion